OpenStudy (anonymous):

Regan is trying to find the equation of a quadratic that has a focus of (−2, 5) and a directrix of y = 13. Describe to Regan your preferred method for deriving the equation. Make sure you use Regan's situation as a model to help her understand.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

a quadratic is a parabola You should first find the vertex, which is halfway between the focus and the directrix what do you get ?

OpenStudy (anonymous):

√(x2--2)^2+(y1-5)^2= √(y2-13)^2

OpenStudy (phi):

what is the vertex ?

OpenStudy (anonymous):

I get this when I simplify it but I don't know the vertex... X^2+(y1-5)^2+2=y2-13

OpenStudy (phi):